### v2

# Kempe equivalence and quadratic toric rings

Abstract.
Perfectly contractile graphs form a typical class of perfect graphs. In particular, all $k$-colorings of a perfectly contractile graph are Kempe equivalent. Everett and Reed conjectured that a graph is perfectly contractile if and only if it contains no odd holes, no antiholes and no odd prisms. On the other hand the authors and Shibata conjectured that a perfect graph is perfectly contractile if and only if its toric ring, which is called the stable set ring, is quadratic. In the present paper, we characterize when the stable set ring of a (not necessarily perfect) graph is quadratic by using Kempe equivalence. As applications of this characterization, we can claim that if Everett and Reed conjecture is true, then the conjecture of the authors and Shibata is also true. Moreover, we can show that for several important classes of perfectly contractile graphs, the stable set rings are quadratic.
Key words and phrases:
Kempe equivalence, stable set ideal, toric ideal, quadratic generated, perfect graph, perfectly contractile graph
2020 Mathematics Subject Classification:
13P10, 05C17
1. Introduction
A graph $G$ is called perfect if every induced subgraph $H$ of $G$ satisfies $\chi(H)=\omega(H)$, where $\chi(H)$ is the chromatic number of $H$ and $\omega(H)$ is the maximum cardinality of cliques of $H$. Perfect graphs were introduced by Berge in [1]. A hole is an induced cycle of length $\geq 5$ and an antihole is the complement of a hole. In [3], Chudnovsky, Robertson, Seymour and Thomas showed that a graph is perfect if and only if it has no odd holes and no odd antiholes. This result is called the strong perfect graph theorem.
On the other hand, Bertschi introduced a hereditary class of perfect graphs in [2]. An even pair in a graph $G$ is a pair of non-adjacent vertices of $G$ such that the length of all induced paths between them is even. Contracting a pair of vertices $\{x,y\}$ in a graph $G$ means removing $x$ and $y$ and adding a new vertex $z$ with edges to every neighborhood of $x$ or $y$. A graph $G$ is called even-contractile if there exists a sequence $G_{0},\ldots,G_{k}$ of graphs satisfying the following:
  1. (i)
    $G=G_{0}$;
  2. (ii)
    each $G_{i}$ is obtained from $G_{i-1}$ by contracting an even pair of $G_{i-1}$;
  3. (iii)
    $G_{k}$ is a complete graph.
A graph $G$ is called perfectly contractile if every induced subgraph of $G$ is even-contractile. Every perfectly contractile graph is perfect, and in particular, each of Meyniel graphs, perfectly orderable graphs and clique separable graphs is perfectly contractile [2]. In addition, every weakly chordal graph is perfectly contractile [6]. Everett and Reed conjectured a forbidden graph characterization of perfectly contractile graphs. An odd (resp. even) prism is a graph consisting of two disjoint triangles with three disjoint induced paths of odd (resp. even) length between them.
Conjecture 1.1 ([5]).
A graph is perfectly contractile if and only if it contains no odd holes, no antiholes and no odd prisms as induced subgraphs.
In [9], the sufficient condition of Conjecture 1.1 was shown. Moreover, Conjecture 1.1 holds for planar graphs [9], dart-free graphs [8], and even prism-free graphs [10]. However, Conjecture 1.1 is still open.
The authors and Shibata extended Conjecture 1.1 to a conjecture on a ring theoretical characterization of perfectly contractile graphs.
Conjecture 1.2 ([13, Conjecture 0.2]).
Let $G$ be a perfect graph. Then the following conditions are equivalent:
  1. (1)
    $G$ is perfectly contractile;
  2. (2)
    $G$ contains no odd holes, no antiholes and no odd prisms as induced subgraphs;
  3. (3)
    the stable set ring $K[G]$ (see Section 2 for the definition) is quadratic.
In [13], (3) $\Rightarrow$ (2) of Conjecture 1.2 was shown. Moreover, if $G$ is one of Meyniel graphs, perfectly orderable graphs and clique separable graphs, then $K[G]$ is quadratic [13].
In the present paper, we characterize when the stable set ring of a (not necessarily perfect) graph is quadratic by using Kempe equivalence. A $k$-coloring $f$ of a graph $G$ on the vertex set $[n]:=\{1,2,\ldots,n\}$ is a map from $[n]$ to $[k]$ such that $f(i)\neq f(j)$ for all $\{i,j\}\in E(G)$. Given a $k$-coloring $f$ of $G$, and $1\leq i<j\leq k$, let $H$ be a connected component of the induced subgraph of $G$ on the vertex set $f^{-1}(i)\cup f^{-1}(j)$. Then we can obtain a new $k$-coloring $g$ of $G$ by setting
$g(x)=\begin{cases}f(x)&x\notin H,\\ i&x\in H\mbox{ and }f(x)=j,\\ j&x\in H\mbox{ and }f(x)=i.\\ \end{cases}$
We say that $g$ is obtained from $f$ by a Kempe switching. Two $k$-colorings $f$ and $g$ are called Kempe equivalent if there exists a sequence $f_{0},f_{1},\ldots,f_{s}$ of $k$-colorings of $G$ such that $f_{0}=f$, $f_{s}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching. See [11] for an overview of Kempe equivalence.
Our first main result gives a system of generators of the defining ideal of $K[G]$ by using $k$-colorings (Theorem 3.3). Moreover, our second main result characterizes quadratic stable set ring by using Kempe equivalence. In fact,
Theorem 1.3.
Let $G$ be a graph on the vertex set $[n]$. Then $K[G]$ is quadratic if and only if, for every replication graph $H$ of an arbitrary induced subgraph of $G$ and every $k\geq\chi(H)$, all $k$-colorings of $H$ are Kempe equivalent (see Section 3 for the definition of replication graphs).
As an application of this theorem, we can claim that Conjecture 1.1 implies Conjecture 1.2.
Theorem 1.4.
If Conjecture 1.1 is true, so is Conjecture 1.2.
Moreover, as an application, it is easy to show that for several important classes of perfectly contractile graphs, the stable set rings are quadratic. In fact,
Theorem 1.5.
For any graph $G$ belonging to one of the following classes of perfectly contractile graphs, $K[G]$ is quadratic.
  • (a)
    dart-free graphs with no odd holes, no antiholes, and no odd prisms;
  • (b)
    even prism-free graphs with no odd holes, no antiholes, and no odd prisms;
  • (c)
    weakly chordal graphs;
  • (d)
    Meyniel graphs [13];
  • (e)
    perfectly orderable graphs [13].
Finally, combining Conjecture 1.2 and our characterization of quadratic stable set rings, we give a new conjecture of a graph-theoretical characterization of perfectly contractile graphs (Conjecture 4.3).
The present paper is organized as follows: First, in Section 2, we define stable set rings. We then, in Section 3, discuss relations between Kempe equivalence and stable set rings. In particular, we prove Theorem 1.3. Finally, in Section 4, we consider applications of Theorem 1.3 and prove Theorems 1.4 and 1.5.
2. Stable set rings
In this section, we define stable set rings. Let $G$ be a graph on the vertex set $[n]$ with the edge set $E(G)$. Given a subset $S\subset[n]$, let $G[S]$ denote the induced subgraph of $G$ on the vertex set $S$. A subset $S\subset[n]$ is called a stable set (or an independent set) of $G$ if $\{i,j\}\notin E(G)$ for all $i,j\in S$ with $i\neq j$. Namely, a subset $S\subset[n]$ is stable if and only if $G[S]$ is an empty graph. In particular, the empty set $\emptyset$ and any singleton $\{i\}$ with $i\in[n]$ are stable. Denote $S(G)=\{S_{1},\ldots,S_{m}\}$ the set of all stable sets of $G$. Given a subset $S\subset[n]$, we associate the $(0,1)$-vector $\rho(S)=\sum_{j\in S}{\mathbf{e}}_{j}$. Here ${\mathbf{e}}_{j}$ is the $j$th unit coordinate vector in ${\mathbb{R}}^{n}$. For example, $\rho(\emptyset)=(0,\ldots,0)\in{\mathbb{R}}^{n}$. Let $K[{\mathbf{t}},s]:=K[t_{1},\ldots,t_{n},s]$ be the polynomial ring in $n+1$ variables over a field $K$. Given a nonnegative integer vector ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in{\mathbb{Z}}_{\geq 0}^{n}$, we write ${\mathbf{t}}^{{\mathbf{a}}}:=t_{1}^{a_{1}}t_{2}^{a_{2}}\cdots t_{n}^{a_{n}}\in K[{\mathbf{t}},s]$. The stable set ring of $G$ is
$K[G]:=K[{\mathbf{t}}^{\rho(S_{1})}s,\ldots,{\mathbf{t}}^{\rho(S_{m})}s]\subset K[{\mathbf{t}},s].$
We regard $K[G]$ as a homogeneous algebra by setting each $\deg({\mathbf{t}}^{\rho(S_{i})}s)=1$. Note that $K[G]$ is a toric ring. Let $K[{\mathbf{x}}]:=K[x_{1},\ldots,x_{m}]$ denote the polynomial ring in $m$ variables over $K$ with each $\deg(x_{i})=1$. The stable set ideal of $G$ is the kernel of the surjective homomorphism $\pi:K[{\mathbf{x}}]\to K[G]$ defined by $\pi(x_{i})={\mathbf{t}}^{\rho(S_{i})}s$ for $1\leq i\leq m$. Note that $I_{G}$ is a toric ideal, and generated by homogeneous binomials. The toric ring $K[G]$ is called quadratic if $I_{G}$ is generated by quadratic binomials. We say that “$I_{G}$ is generated by quadratic binomials” even if $I_{G}=\{0\}$ (or equivalently, $G$ is complete). It is easy to see that a homogeneous binomial $x_{i_{1}}\cdots x_{i_{r}}-x_{j_{1}}\cdots x_{j_{r}}\in K[{\mathbf{x}}]$ belongs to $I_{G}$ if and only if $\bigcup_{\ell=1}^{r}S_{i_{\ell}}=\bigcup_{\ell=1}^{r}S_{j_{\ell}}$ as multisets. See, e.g., [7] for details on toric rings and toric ideals.
3. Kempe equivalence and quadratic stable set rings
In this section, we characterize when the stable set ring $K[G]$ of a (not necessarily perfect) graph is quadratic. In particular, we prove Theorem 1.3. First, we discuss a relationship between monomials of degree $k$ in $K[{\mathbf{x}}]$ and $k$-colorings.
Given a graph $G$ on the vertex set $[n]$, and ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in{\mathbb{Z}}_{\geq 0}^{n}$, let $G_{\mathbf{a}}$ be the graph obtained from $G$ by replacing each vertex $i\in[n]$ with a complete graph $G^{(i)}$ of $a_{i}$ vertices (if $a_{i}=0$, then just delete the vertex $i$), and joining all vertices $x\in G^{(i)}$ and $y\in G^{(j)}$ such that $\{i,j\}$ is an edge of $G$. In particular, if ${\mathbf{a}}=(1,\ldots,1)$, then $G_{\mathbf{a}}=G$. If ${\mathbf{a}}={\bf 0}$, then $G_{\mathbf{a}}$ is the null graph (a graph without vertices). In addition, if ${\mathbf{a}}$ is a $(0,1)$-vector, then $G_{\mathbf{a}}$ is an induced subgraph of $G$. If ${\mathbf{a}}$ is a positive vector, then $G_{\mathbf{a}}$ is called a replication graph of $G$. In general, $G_{\mathbf{a}}$ is a replication graph of an induced subgraph of $G$.
Lemma 3.1.
Let $G$ be a graph on the vertex set $[n]$ and let $f$ be a $k$-coloring of $G_{{\mathbf{a}}}$ where ${\mathbf{a}}\in{\mathbb{Z}}^{n}_{\geq 0}$. Then for each $\ell=1,2,\ldots,k$, the set $S=\{j\in[n]:G^{(j)}\cap f^{-1}(\ell)\neq\emptyset\}$ is a stable sets of $G$.
Proof.
Suppose that $\{j_{1},j_{2}\}$ is an edge of $G$ for some $j_{1},j_{2}\in S$. Since $G^{(j_{1})}\cap f^{-1}(\ell)\neq\emptyset$ and $G^{(j_{2})}\cap f^{-1}(\ell)\neq\emptyset$ hold, there exist $v_{1}\in G^{(j_{1})}\cap f^{-1}(\ell)$ and $v_{2}\in G^{(j_{2})}\cap f^{-1}(\ell)$. Since $\{j_{1},j_{2}\}$ is an edge of $G$, $\{v_{1},v_{2}\}$ is an edge of $G_{\mathbf{a}}$. However, since $v_{1},v_{2}\in f^{-1}(\ell)$, this contradicts that $f$ is a $k$-coloring of $G_{\mathbf{a}}$. Thus $S$ is a stable set of $G$. ∎
Let $G$ be a graph on the vertex set $[n]$ with stable sets $S(G)=\{S_{1},\ldots,S_{m}\}$ and let $f$ be a $k$-coloring of $G_{\mathbf{a}}$ where ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$. From Lemma 3.1 we can associate $f$ with a monomial
(1) $\varphi_{k}(f):=x_{i_{1}}\cdots x_{i_{k}}\in K[{\mathbf{x}}],$
where $S_{i_{\ell}}=\{j\in[n]:G^{(j)}\cap f^{-1}(\ell)\neq\emptyset\}$ for $\ell=1,2,\ldots,k$.
Lemma 3.2.
Let $G$ be a graph on the vertex set $[n]$. Then, for any positive integer $k$, $\varphi_{k}$ given in (1) is a surjective map from $\{f:f\mbox{ is a }k\mbox{-coloring of }G_{\mathbf{a}}\mbox{ for some }{\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}\}$ to the set of all monomials of degree $k$ in $K[{\mathbf{x}}]$. Moreover, $\varphi_{k}(f)=\varphi_{k}(g)$ if and only if $f$ is obtained from $g$ by either exchanging colors or exchanging the coloring of vertices in each clique $G^{(j)}$ of $G_{\mathbf{a}}$.
Proof.
Let $M=x_{i_{1}}\cdots x_{i_{k}}$ be a monomial in $K[{\mathbf{x}}]$. Let ${\mathbf{a}}=(a_{1},\ldots,a_{n})$ be the integer vector defined by $a_{p}=|\{\ell:p\in S_{i_{\ell}}\}|$. Then $S_{i_{1}},\ldots,S_{i_{k}}$ defines a $k$-coloring $f$ of $G_{\mathbf{a}}$ as follows: if $S_{i_{\ell}}$ contains $j$, then let $f(x)=\ell$ for one of $x\in G^{(j)}$ which is not colored yet. Then $\varphi_{k}(f)=M$. Thus $\varphi_{k}$ is surjective. This coloring is unique up to the choices of a vertex from $G^{(j)}$. ∎
Since a homogeneous binomial $x_{i_{1}}\cdots x_{i_{k}}-x_{j_{1}}\cdots x_{j_{k}}\in K[{\mathbf{x}}]$ belongs to $I_{G}$ if and only if $\bigcup_{\ell=1}^{k}S_{i_{\ell}}=\bigcup_{\ell=1}^{k}S_{j_{\ell}}$ as multisets (i.e., |$|\{\ell:p\in S_{i_{\ell}}\}|=|\{\ell:p\in S_{j_{\ell}}\}|$ for any $1\leq p\leq n$), we can give a system of generators of $I_{G}$ as follows.
Theorem 3.3.
Let $F=x_{i_{1}}\cdots x_{i_{k}}-x_{j_{1}}\cdots x_{j_{k}}\in K[{\mathbf{x}}]$ be a homogeneous binomial. Then $F\in I_{G}$ if and only if there exists ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$ such that $\varphi_{k}(f)=x_{i_{1}}\cdots x_{i_{k}}$ and $\varphi_{k}(g)=x_{j_{1}}\cdots x_{j_{k}}$ for some $k$-colorings $f$ and $g$ of $G_{\mathbf{a}}$. In particular, one has
$I_{G}=\langle\varphi_{k}(f)-\varphi_{k}(g):\mbox{$f$ and $g$ are k-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in{\mathbb{Z}}^{d}_{\geq 0}$ and $k\geq\chi(G_{{\mathbf{a}}})$}\rangle.$
Now, we can prove Theorem 1.3.
Proof of Theorem 1.3.
(only if) Suppose that $K[G]$ is quadratic. Let $f$ and $g$ be $k$-colorings of $G_{\mathbf{a}}$ with ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$ which are not Kempe equivalent. From Theorem 3.3, the binomial
$F:=\varphi_{k}(f)-\varphi_{k}(g)=x_{i_{1}}\cdots x_{i_{k}}-x_{j_{1}}\cdots x_{j_{k}}$
belongs to $I_{G}$. By the hypothesis, $F$ is generated by quadratic binomials in $I_{G}$. Then the theory of binomial ideals [7, Lemma 3.8] guarantees that there exists an expression
(2) $F=\sum_{r=1}^{s}{\bf x}^{{\mathbf{w}}_{r}}(x_{p_{r}}x_{q_{r}}-x_{p_{r}^{\prime}}x_{q_{r}^{\prime}}),$
where each $x_{p_{r}}x_{q_{r}}-x_{p_{r}^{\prime}}x_{q_{r}^{\prime}}$ belongs to $I_{G}$. Suppose that $s$ is minimal among all pairs of $k$-colorings of $G_{\mathbf{a}}$ which are not Kempe equivalent. Since $x_{i_{1}}\cdots x_{i_{k}}$ must appear in the right-hand side of (2), we may assume that $x_{i_{1}}\cdots x_{i_{k}}={\bf x}^{{\mathbf{w}}_{1}}x_{p_{1}}x_{q_{1}}$. Then there exist $\ell$ and $\ell^{\prime}$ such that $x_{i_{\ell}}x_{i_{\ell^{\prime}}}=x_{p_{1}}x_{q_{1}}$, say, $x_{i_{1}}x_{i_{2}}=x_{p_{1}}x_{q_{1}}$. Note that $\varphi_{2}(f_{1})=x_{i_{1}}x_{i_{2}}$, where $f_{1}$ is the restriction of $f$ to $G^{\prime}=G_{\mathbf{a}}[f^{-1}(1)\cup f^{-1}(2)]$. Moreover since $x_{p_{1}}x_{q_{1}}-x_{p_{1}^{\prime}}x_{q_{1}^{\prime}}=x_{i_{1}}x_{i_{2}}-x_{p_{1}^{\prime}}x_{q_{1}^{\prime}}$ belongs to $I_{G}$, $\varphi_{2}(f_{2})=x_{p_{1}^{\prime}}x_{q_{1}^{\prime}}$ for some 2-coloring $f_{2}$ of the graph $G^{\prime}$. Since $G^{\prime}$ has a 2-coloring, it is a bipartite graph. Then $f_{1}$ and $f_{2}$ are Kempe equivalent (See, e.g., [11, Proposition 2.1 (a)]). Let $f^{\prime}$ be a coloring of $G_{\mathbf{a}}$ defined by
$f^{\prime}(x)=\begin{cases}f_{2}(x)&f(x)\in\{1,2\},\\ f(x)&\mbox{otherwise.}\end{cases}$
It then follows that $f$ and $f^{\prime}$ are Kempe equivalent. Moreover we have
$x_{i_{1}}\cdots x_{i_{k}}-{\bf x}^{{\mathbf{w}}_{1}}(x_{p_{1}}x_{q_{1}}-x_{p_{1}^{\prime}}x_{q_{1}^{\prime}})=x_{p_{1}^{\prime}}x_{p_{2}^{\prime}}x_{i_{3}}\cdots x_{i_{k}}=\varphi_{k}(f^{\prime}).$
Since
$\varphi_{k}(f^{\prime})-\varphi_{k}(g)=x_{p_{1}^{\prime}}x_{p_{2}^{\prime}}x_{i_{3}}\cdots x_{i_{k}}-x_{j_{1}}\cdots x_{j_{k}}=\sum_{r=2}^{s}{\bf x}^{{\mathbf{w}}_{r}}(x_{p_{r}}x_{q_{r}}-x_{p_{r}^{\prime}}x_{q_{r}^{\prime}}),$
$f^{\prime}$ and $g$ are Kempe equivalent by the hypothesis on $s$. Since $f$ and $f^{\prime}$ are Kempe equivalent, this contradicts that $f$ and $g$ are not Kempe equivalent.
(if) Suppose that for every ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$, all $k$-colorings of $G_{\mathbf{a}}$ are Kempe equivalent. Let $F=x_{i_{1}}\cdots x_{i_{k}}-x_{j_{1}}\cdots x_{j_{k}}\in I_{G}$ be a homogeneous binomial of degree $k\geq 3$. Suppose that $F$ is not generated by quadratic binomials in $I_{G}$.
From Theorem 3.3, there exist $k$-colorings $f$ and $g$ of $G_{\mathbf{a}}$ such that $\varphi_{k}(f)=x_{i_{1}}\cdots x_{i_{k}}$ and $\varphi_{k}(g)=x_{j_{1}}\cdots x_{j_{k}}$, where ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in{\mathbb{Z}}_{\geq 0}^{n}$ is given by $a_{p}=|\{\ell:p\in S_{i_{\ell}}\}|\ (=|\{\ell:p\in S_{j_{\ell}}\}|)$. By the hypothesis, $f$ and $g$ are Kempe equivalent. Let $f_{0},f_{1},\ldots,f_{s}$ be a sequence of $k$-colorings of $G_{\mathbf{a}}$ such that $f_{0}=f$, $f_{s}=g$, and $f_{i}$ is obtained from $f_{i-1}$ by a Kempe switching. Suppose that $s$ is minimal among all binomials which are not generated by quadratic binomials in $I_{G}$.
Suppose that the Kempe switching from $f$ to $f_{1}$ is obtained by a connected component $H$ of the induced subgraph $G_{\mathbf{a}}[f^{-1}(\mu)\cup f^{-1}(\eta)]$ by setting
$f_{1}(x)=\begin{cases}f(x)&x\notin H,\\ \mu&x\in H\mbox{ and }f(x)=\eta,\\ \eta&x\in H\mbox{ and }f(x)=\mu.\\ \end{cases}$
Let $f|_{G^{\prime}}$ and $f_{1}|_{G^{\prime}}$ be the restrictions of $f$ and $f_{1}$ to $G^{\prime}:=G_{\mathbf{a}}[f^{-1}(\mu)\cup f^{-1}(\eta)]$, respectively. Since $f|_{G^{\prime}}$ and $f_{1}|_{G^{\prime}}$ are 2-colorings of ${G^{\prime}}$, there exists a quadratic binomial $x_{i_{\mu}}x_{i_{\eta}}-x_{\sigma}x_{\tau}\in I_{G}$ where $S_{\sigma}=\{j\in[n]:G^{(j)}\cap f_{1}^{-1}(\mu)\neq\emptyset\}$ and $S_{\tau}=\{j\in[n]:G^{(j)}\cap f_{1}^{-1}(\eta)\neq\emptyset\}$. Then
$F=\frac{x_{i_{1}}\cdots x_{i_{k}}}{x_{i_{\mu}}x_{i_{\eta}}}(x_{i_{\mu}}x_{i_{\eta}}-x_{\sigma}x_{\tau})+F^{\prime},$
where
$F^{\prime}=\frac{x_{i_{1}}\cdots x_{i_{k}}}{x_{i_{\mu}}x_{i_{\eta}}}x_{\sigma}x_{\tau}-x_{j_{1}}\cdots x_{j_{k}}\in I_{G}.$
Moreover $F^{\prime}=\varphi_{k}(f_{1})-\varphi_{k}(g)$ for the $k$-colorings $f_{1}$ and $g$ of $G_{\mathbf{a}}$. By the hypothesis on $s$, $F^{\prime}$ is generated by quadratic binomials $F_{1},\ldots,F_{t}$ in $I_{G}$. Hence $F$ is generated by quadratic binomials $x_{i_{\mu}}x_{i_{\eta}}-x_{\sigma}x_{\tau},F_{1},\ldots,F_{t}$ in $I_{G}$, a contradiction. ∎
By using Theorem 1.3, we give an example of even-contractile graph $G$ such that $K[G]$ is not quadratic.
Example 3.4.
Let $G$ be the following graph:
$G:$1111222233334444555566667777
Then $G$ is even-contractile. Indeed, in the following graphs, each pair of the black vertices is an even pair and each $G_{i}$ is obtained from $G_{i-1}$ by contracting the even pair of $G_{i-1}$.
$G_{0}:$$G_{1}:$$G_{2}:$$G_{3}:$
It then follows that for every $k\geq\chi(G)=4$, all $k$-colorings of $G$ are Kempe equivalent (see Proposition 4.1). On the other hand, for ${\mathbf{a}}=(1,1,1,1,1,1,0)$, $G_{\mathbf{a}}$ is an odd prism which has 3-colorings $f$ and $g$ defined by
$f(i)=\begin{cases}1&\mbox{if $i\in\{1,5\}$}\\ 2&\mbox{if $i\in\{2,6\}$}\\ 3&\mbox{if $i\in\{3,4\}$}\end{cases}\mbox{ and }g(i)=\begin{cases}1&\mbox{if $i\in\{1,6\}$}\\ 2&\mbox{if $i\in\{2,4\}$}\\ 3&\mbox{if $i\in\{3,5\}$}.\end{cases}$
Clearly, $f$ and $g$ are not Kempe equivalent (as observed in [15]). Hence from Theorem 1.3, $K[G]$ is not quadratic. In fact, setting
$S_{1}=\{1,5\},S_{2}=\{2,6\},S_{3}=\{3,4\},S_{4}=\{1,6\},S_{5}=\{2,4\},S_{6}=\{3,5\},$
each $S_{i}$ is a stable set of $G$ and $F=x_{1}x_{2}x_{3}-x_{4}x_{5}x_{6}$ is a binomial in $I_{G}$ which is not generated by quadratic binomials in $I_{G}$.
Remark 3.5.
It is known that $G$ is perfect if and only if the initial ideal of $I_{G}$ is generated by squarefree monomials with respect to any reverse lexicographic order [12]. In such a case, $I_{G}$ is generated by the binomials of degree $\leq n+1$ from [14, Proposition 13.15 (ii)]. Hence, if $G$ is perfect, then we can replace “$k\geq\chi(H)$” by “$\chi(H)\leq k\leq n+1$” in Theorem 1.3.
4. Applications to perfectly contractile graphs
In this section, we consider applications of Theorem 1.3 and prove Theorems 1.4 and 1.5. The following result is a motivation for Bertschi to define perfectly contractile graphs.
Proposition 4.1 ([2]).
Let $G$ be an even-contractile graph. Then $\chi(G)=\omega(G)$ and for every $k\geq\chi(G)$, all $k$-colorings are Kempe equivalent.
By using this proposition and Theorem 1.3, we have the following.
Corollary 4.2.
Let $\mathcal{C}$ be a class of some perfectly contractile graphs. Suppose that, for any $G\in\mathcal{C}$ and ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$, the graph $G_{\mathbf{a}}$ belongs to $\mathcal{C}$. Then $K[G]$ is quadratic for every $G\in\mathcal{C}$.
Proof.
Let $G\in\mathcal{C}$ and ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$. Since $G_{\mathbf{a}}\in\mathcal{C}$ is perfectly contractile, all $k$-colorings of $G_{\mathbf{a}}$ are Kempe equivalent by Proposition 4.1. From Theorem 1.3, $K[G]$ is quadratic. ∎
A pair of adjacent vertices $\{x,y\}$ of $G$ is called an adjacent twin in $G$ if $N(x)\setminus\{y\}=N(y)\setminus\{x\}$, where $N(x)$ is the set of all neighbours of $x$ in $G$. By using Corollary  4.2, we can prove Theorem 1.4.
Proof of Theorem 1.4.
Suppose that Conjecture 1.1 is true, that is, the class $\mathcal{A}$ of graphs that contain no odd holes, no antiholes, and no odd prisms coincides with the set of all perfectly contractile graphs. Since (3) $\Rightarrow$ (2) is known, it is enough to show (2) $\Rightarrow$ (3).
Let $G\in\mathcal{A}$. From Corollary 4.2, it is enough to show that, for any ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$, we have $G_{\mathbf{a}}\in\mathcal{A}$. Since the class $\mathcal{A}$ is hereditary, we may assume that ${\mathbf{a}}$ is a positive vector. Note that, if $x,y\in G^{(i)}$ with $x\neq y$, then $\{x,y\}$ is an adjacent twin. Since odd holes, antiholes, and odd prisms have no adjacent twins, $G_{\mathbf{a}}$ has one of such graphs as an induced subgraph if and only if so does $G$. Hence we have $G_{\mathbf{a}}\in\mathcal{A}$. Thus (2) $\Rightarrow$ (3) holds. ∎
We apply Corollary 4.2 to several important classes of perfectly contractile graphs to prove Theorem 1.5. A dart is a graph on the vertex set $\{1,2,3,4,5\}$ and the edge set
$\{\{1,2\},\{2,3\},\{1,5\},\{2,5\},\{3,5\},\{4,5\}\}.$
A graph is called weakly chordal if it has no holes and no antiholes. A graph is called Mayniel if any odd cycle of length $\geq 5$ has at least two chords. Let $G$ be a graph on the vertex set $[n]$. An ordering $(i_{1},\ldots,i_{n})$ of $[n]$ is called perfect if, for any induced (ordered) subgraph $H$ of $G$, the number of colors used by Greedy Coloring on $H$ coincides with $\chi(H)$. It is known [4, Theorem 1] that a vertex ordering $<$ of a graph $G$ is perfect if and only if $G$ contains no induced $P_{4}$ $abcd$ such that $a<b$ and $d<c$. A graph is called perfectly orderable if there exists a perfect ordering $(v_{1},\dots,v_{n})$ of the vertex set of $G$.
Proof of Theorem 1.5.
From Corollary 4.2, it is enough to show that, for any ${\mathbf{a}}\in{\mathbb{Z}}_{\geq 0}^{n}$, if $G$ satisfies one of (a) – (e), then so does $G_{\mathbf{a}}$. Since all of the classes are hereditary, we may assume that ${\mathbf{a}}$ is positive.
A dart, holes, antiholes, odd/even prisms, and odd cycles of length $\geq 5$ having exactly one chord are forbidden graphs for the classes (a), (b), (c) and (d), and have no adjacent twins. Thus $G_{\mathbf{a}}$ has one of such graphs as an induced subgraph if and only if so does $G$.
Let $G$ be a perfectly orderable graph with a perfect ordering $(i_{1},\ldots,i_{n})$. Note that induced $P_{4}$ has no adjacent twins. Hence any vertex ordering of $G_{\mathbf{a}}$ such that $x<y$ whenever $x\in G^{(i)}$, $y\in G^{(j)}$ with $i<j$ is a perfect ordering of $G_{\mathbf{a}}$. Thus $G_{\mathbf{a}}$ is perfectly orderable. ∎
Finally, from Conjecture 1.2 and Theorem 1.3, we conjecture a new characterization of perfectly contractile graphs.
Conjecture 4.3.
Let $G$ be a perfect graph. Then $G$ is perfectly contractile if and only if for every replication graph $H$ of an arbitrary induced subgraph of $G$ and every $k\geq\chi(H)$, all $k$-colorings of $H$ are Kempe equivalent.
Acknowledgment
The first author was partially supported by JSPS KAKENHI 18H01134 and the second author was partially supported by JSPS KAKENHI 22K13890.